OpenStudy (anonymous):

What is the factored form of the following expressions? 17. x2 – 10xy + 24y2

OpenStudy (hoblos):

x2 – 10xy + 24y2= x2 – 10xy + 25y2 -y2 = (x-5y)^2 -y2 = (x-5y-y)(x-5y+y) = (x-6y)(x-4y)